# An Adjoint Formulation of Energetic Particle Confinement

### v1

@@ III.2 Jonta: A Particle Based Drift Kinetic Solver
Particle based drift kinetic simulations of fast ion orbits is computationally intensive White and Chance (1984); Hirvijoki et al. (2014), which requires moderate to substantial resources and efficient software. While this has predominately been done with Fortran or C++ libraries, with GPU acceleration facilitated by libraries such as Kokkos Trott et al. (2021), maintaining the support of these solvers can be cumbersome, due to its dependency on external backend libraries to effectively utilize high performance computing (HPC) hardware. Perhaps most importantly, given that PINNs are trained on GPUs and libraries with a large user base, this motives the development of a traditional drift kinetic solver that runs on GPUs and utilizes the same libraries that are used by PINNs, facilitating the operation of the broader framework that contains an interplay between the PINN and a drift kinetic solver.
With the aforementioned discussion, we present Just anOther fuNcTionAl pusher (JONTA), which is built on the PyTorch Paszke et al. (2019) and JAX Bradbury et al. (2018) libraries that run on GPU accelerated hardware. For the present work, we will utilize the JAX backend, and employ a four stage Runge Kutta (RK4) integration scheme, which is part of a broader suite containing differentiable integrator methods in the Optax library opt (2020). Guiding center equations are evolved with RK4, and collisions are implemented with a Monte-Carlo operator described below. Upon acceptance of the manuscript, the code will be uploaded to the public repository, where additional implementations can also be found for evolving charged particles in other fusion scenarios, such as runaway electrons in tokamak disruptions Breizman et al. (2019).
@@
where we have chosen cylindrical coordinates $\left(R,Z\right)$, $X\equiv R-R_{0}$, and make the transformation from $\left(\mathbf{x},v_{\|},\mu\right)$ to $\left(\mathbf{x},\xi,v\right)$. Assuming an axisymmetric tokamak geometry with circular geometry and noting the relations $\xi^{2}=v^{2}_{\|}/\left(v^{2}_{\|}+2\mu B\right)$, and $v^{2}=v^{2}_{\|}+2\mu B$, Eq. (14) can be written as
@@
Here, we normalized time to $a/v_{Ti}$, where $a$ is the plasma minor radius taken to be equal to $r_{wall}$, speed has been normalized to $v_{Ti}$, space to $a$, and we have defined $\rho^{*}_{i}=\rho_{i}/a$, $\rho_{i}=v_{Ti}/\omega^{(0)}_{c}$, $\omega^{(0)}_{c}=ZeB_{0}/m_{i}$, $B_{0}$ is the magnetic field on axis, and $q$ is the safety factor. For convenience we have neglected the electric field, and assumed the limit of a small inverse aspect ratio and low-$\beta$ plasma. Specifically, the magnetic field was taken to have the form:
@@
The collision operator is taken to be Lorentz operator [Eq. (2)], where the deflection frequency in the limit of $v\gg v_{Ti}$ can be expressed as:
$\frac{a\nu_{D}}{v_{Ti}}=\left(a\frac{\hat{\nu}_{si}}{v_{Ti}}\right)\frac{v^{3}_{Ti}}{v^{3}}=\pi\left(n_{i}r^{2}_{e}a\right)Z^{2}_{s}Z^{2}_{i}\left(\frac{m_{e}c^{2}}{T_{i}}\right)^{2}\ln\Lambda\frac{v^{3}_{Ti}}{v^{3}},$
@@
where $\Delta t$ is the collisional time step, which is taken to be the same as the collisionless time step. While the Lorentz collision operator does not account for the slowing down of energetic ions, a critical aspect of describing energetic particle evolution, it does drive cross field transport. This will thus allow us to identify energetic ions that are lost rapidly, versus those that will be much better confined. The addition of a more complete set of physical processes including 3D magnetic fields, energetic particle modes, and collisional slowing down will be pursued in future work.
@@ III.3 Example Ion Orbits and Coordinate Conventions
Figure 2: Circular flux surface geometry used in this work and example collisionless ion orbits. Panel (a) shows co-current ions with a passing ion with initial pitch $\xi=0.8$ (red curve), and a trapped ion with initial pitch $\xi=0.3$ (green curve). Panel (b) shows counter-current ions with a passing ion with initial pitch $\xi=-0.8$ (red curve), and a trapped ion with initial pitch $\xi=-0.3$ (green curve). Flux surface contours are shown in blue, with the example orbits shown in red. The black marker indicates the initial position of the ions, and arrows indicate their direction. The grad-B drift points downward in our example geometry. A deuterium ion with $20\;\text{keV}$, a tokamak with a minor radius of $0.5\;\text{m}$, inverse aspect ratio $a/R_{0}=1/3$, constant $q$-profile of $q=2$, and an on-axis magnetic field strength of $B_{0}=2\;\text{T}$ was assumed.
@@
Before discussing solutions to the adjoint drift kinetic equation, it will be useful to describe the magnetic geometry employed. A plot of the circular flux surfaces assumed together with four example ion orbits are shown in Fig. 2. Here, the blue contours indicate magnetic flux surfaces, and the red and green curves indicate example passing and trapped ion orbits, respectively. The toroidal magnetic field and current are assumed to come out of the page, with the poloidal magnetic field in the counter-clockwise direction. Passing ions with a positive pitch (co-current), will thus rotate in the counter-clockwise direction when their motion is projected onto the poloidal plane as they propagate along a magnetic field line, with a grad-B drift vertically downward. Ions with a negative pitch (counter-current) will move in a clockwise sense when moving along field lines, but will still have a vertically downward grad-B drift. From Fig. 2, it is apparent that co-current ions [Fig. 2(a)] born on the weak field side will tend to be better confined compared to counter-current ions [Fig. 2(b)] born at the same location, due to the magnetic drifts taking the counter-current ions closer to the spatial boundary. Furthermore, a counter-current propagating passing ion when scattered into the trapped region will move onto an orbit that takes it closer to the plasma edge, as evident from Fig. 2(b), such that these ions will be more susceptible to being lost.
IV Escape Time
@@
Figure 3: Loss histories for 20 keV (panel a.) and 50 keV (panel b.) ions. Solid lines indicate the training loss whereas ‘x’ markers indicate the test loss. Six million training points were used and two million test points.
Figure 4: (a) Escape time for an ion initialized in the counter-current direction with a pitch of $\xi=-0.8$, after different stages of training. The profile of the ion’s escape time is shown in the panels (a), (b) and (c), with the residual indicated in panels (d), (e) and (f). The first column is after the SOAP phase of training (50,000 epochs), the second column is 10,000 into the SSBroyden phase (60,000 net epochs), and the final column is after 50,000 epochs of SSBroyden (100,000 epochs total). The quantity plotted is $\text{log}_{10}\left(1+T\right)$, yielding a $\text{log}_{10}$ scale for large values of $T$, but vanishing when $T=0$. A deuterium ion with $10\;\text{keV}$ was assumed.
@@
This section will describe solutions for the escape time for a fast ion population at 20 and 50 keV utilizing the PINN implementation described in Sec. III.1, along with a comparison to solutions evaluated with JONTA. Loss histories for ions with 20 and 50 keV are shown in Fig. 3. For these cases, we assumed a midsize tokamak with a minor radius of $a=0.5\;\text{m}$, aspect ratio $R_{0}/a=3$, and deuterium fast ions. The background deuterium plasma is assumed to have a temperature of $1\;\text{keV}$, a density of $n_{i}=10^{20}\;\text{m}^{-3}$, and the safety factor is taken to be $q=2$. A representative Coulomb logarithm of $\ln\Lambda=15$ was used in the collision operator. Each of these quantities were taken to be constant across the plasma radius, though spatial profiles could easily be introduced. A fully connected feedforward network with seven layers each with a width of fifty neurons is used. The SOAP optimizer is used for the first 50,000 epochs, with SSBroyden used for the remaining 25,000 epochs. From Fig. 3, it is evident that SOAP drives a slow, but steady decay of the loss, with SSBroyden driving a sharper decline. We note, however, that the SOAP optimizer being employed in this paper is able to effectively use the Blackwell 200 GPU employed in this work, whereas the SciPy implementation of SSBroyden primarily uses a CPU. As a result, the vast majority of the training time is spent during the SSBroyden phase of optimization.
@@
Considering the evolution of the predicted solution during training, Fig. 4(a) shows the predicted escape time for a 20 keV ion after the SOAP optimization phase, 10,000 epochs into the SSBroyden phase [Fig 4(b)], and the final predicted escape time [Fig 4(c)]. A reference solution from JONTA is shown in Fig. 9(d) below for comparison. It is evident that SOAP is able to capture the structure of the solution near the tokamak edge, but fails to recover the long confinement time of ions initially located at small minor radii. This is due to the large discrepancy in time scales between ions that begin on loss orbits, and are thus lost on a transit time scale, and those initially located deep in the plasma interior, which must collisionally diffuse out of the plasma. This latter process takes place on a far longer time scale, roughly $10^{5}$ slower for the present example, which SOAP struggles to recover. After 10,000 epochs with SSBroyden, the training and test loss of the residual and boundary terms drop further, and it is evident from Fig. 4 that the relatively long confinement time of ions initially located in the interior is better captured, though its value is still substantially under predicted. Finally, after 25,000 epochs of SSBroyden, the predicted escape time of ions initially located in the interior slowly increases, though the final value is still below that predicted by the particle-based code JONTA.
@@
Figure 5: Test loss distribution for 20 keV ions. A uniform random distribution was used, where the size of the markers is proportional to the magnitude of the residual.
@@
Turning to the distribution of training losses, six million training points were used with two million uniformly distributed test points applied. The training points were taken to be distributed by a Hammersley sequence. Two dimensional projections of the resulting test error are shown in Fig. 5, with the magnitude of the markers set to be proportional to the magnitude of the residual. Hence, darker regions represent regions with the largest residuals. From Fig. 5, the outer thirty percent of the tokamak contained the largest losses. This follows due to this region containing direct loss orbits, whose detailed form can be difficult to resolve. It is also evident that the outer five percent of the plasma radius has somewhat smaller residuals due to the concentration of training points in this region. Finally, regions with $r>0.99$ and poloidal angles on either the outboard ($\Delta\theta<\theta<\Delta\theta$) or inboard ($\theta>-\pi+\Delta\theta$ and $\theta<\pi-\Delta\theta$) midplanes where set to zero, leading to empty regions in Fig. 5(a). For this example we set $\Delta\theta=0.1$. As described further below, these regions contain discontinuities in the solution arising from changes in boundary conditions between the lower and upper half of the tokamak, which cannot be precisely resolved. By removing training points from this region, this allows the PINN to find a solution that smoothly transitions across these narrow regions.

### v2

@@ III.2 JONTA: A Particle Based Drift Kinetic Solver
Particle-based drift kinetic simulations of fast ion orbits are computationally intensive White and Chance (1984); Hirvijoki et al. (2014) requiring substantial computational resources and efficient software. While this has predominately been done with Fortran or C++ libraries, with GPU acceleration facilitated by libraries such as Kokkos Trott et al. (2021), maintaining the support of these solvers can be cumbersome, due to its dependency on external backend libraries to effectively utilize high performance computing (HPC) hardware. Perhaps most importantly, given that PINNs are trained on GPUs and libraries with a large user base, this motivates the development of a traditional drift kinetic solver that runs on GPUs and utilizes the same libraries that are used by PINNs, facilitating the operation of the broader framework that contains an interplay between the PINN and a drift kinetic solver.
With the aforementioned discussion, we present Just anOther fuNcTionAl pusher (JONTA), which is built on the PyTorch Paszke et al. (2019) and JAX Bradbury et al. (2018) libraries that run on GPU accelerated hardware. For the present work, we will utilize the JAX backend, and employ a four stage Runge Kutta (RK4) integration scheme, which is part of a broader suite containing differentiable integrator methods in the Optax library opt (2020). Guiding-center equations are evolved with RK4, and collisions are implemented with a Monte-Carlo operator described below. The version of JONTA used in this manuscript, together with the PINN training script are available on github.com/cmcdevitt2/DeepPlasma.
@@
where we have chosen cylindrical coordinates $\left(R,Z\right)$, defined $X\equiv R-R_{0}$, and made the transformation from $\left(\mathbf{x},v_{\|},\mu\right)$ to $\left(\mathbf{x},\xi,v\right)$. Assuming an axisymmetric tokamak geometry with circular flux surfaces and noting the relations $\xi^{2}=v^{2}_{\|}/\left(v^{2}_{\|}+2\mu B\right)$ and $v^{2}=v^{2}_{\|}+2\mu B$, Eq. (14) can be written as
@@
Here, we normalized time to $a/v_{Ti}$, where $a$ is the plasma minor radius taken to be equal to $r_{wall}$, speed $v$ has been normalized to $v_{Ti}$, space to $a$, and we have defined $\rho^{*}_{i}=\rho_{i}/a$, $\rho_{i}=v_{Ti}/\omega^{(0)}_{c}$, $\omega^{(0)}_{c}=ZeB_{0}/m_{i}$, $B_{0}$ is the magnetic field on axis, and $q$ is the safety factor. For convenience we have neglected the electric field, and assumed the limit of a small inverse aspect ratio and low-$\beta$ plasma. Specifically, the magnetic field was taken to have the form:
@@
The collision operator is taken to be a Lorentz operator [Eq. (2)], where the deflection frequency in the limit of $v\gg v_{Ti}$ can be expressed as:
$\frac{a\nu_{D}}{v_{Ti}}=\left(a\frac{\hat{\nu}_{si}}{v_{Ti}}\right)\frac{v^{3}_{Ti}}{v^{3}}=\pi\left(n_{i}r^{2}_{e}a\right)Z^{2}_{s}Z^{2}_{i}\left(\frac{m_{e}c^{2}}{T_{i}}\right)^{2}\ln\Lambda\frac{v^{3}_{Ti}}{v^{3}},$ (23)
@@
where $\Delta t$ is the collisional time step, which is taken to be the same as the collisionless time step. While the Lorentz collision operator does not account for the slowing down of energetic ions, a critical aspect of describing energetic particle evolution, pitch-angle scattering will result in collisional cross field transport. The addition of a more complete set of physical processes including 3D magnetic fields, energetic particle modes, and collisional slowing down will be pursued in future work.
Figure 2: (a) Comparison of neoclassical diffusivity computed from JONTA (red ‘x’ markers) and the analytic expression given by Eq. (25) (solid blue curve). (b) Ratio of the analytic expression for the neoclassical diffusivity to JONTA’s prediction. (c) Example fit of squared radial ion displacement versus time (dashed blue curve) with the numerically computed evolution (solid red curve) for particles with an initial radius of $\approx 0.51$. One million total marker particles were used, randomly distributed over pitch and radius. Spatially resolved estimates of transport were made by binning marker particles based on their initial location. The parameters were chosen to be $a=2\;[\text{m}]$, $a/R_{0}=1/6$, $B_{0}=10\;[\text{T}]$, $n_{i}=10^{18}\;[\text{1/m}^{3}]$, $q=2$, $\ln\Lambda=15$, and the ions were assumed to have an energy of $20\;[\text{keV}]$.
To verify our implementation of the Monte Carlo collision operator we compared JONTA predictions of neoclassical transport with an analytic expression for the neoclassical diffusivity. Using the radial coordinate $r=\sqrt{X^{2}+Z^{2}}$, rather than the toroidal flux $\psi_{t}=r^{2}B_{0}/2$, the expression for the neoclassical diffusivity derived in Refs. Rosenbluth et al. (1972); Boozer and Kuo-Petravic (1981) can be written as:
$\frac{D_{neo}}{av_{Ti}}=0.689\sqrt{2\varepsilon}q^{2}\left(\frac{R_{0}}{r}\right)^{2}\left(\frac{m_{s}v^{2}}{2T_{i}}\right)\left(\frac{a\nu_{D}}{v_{Ti}}\right),$ (25)
where $\varepsilon=r/R_{0}$, and $\nu_{D}$ is the pitch-angle scattering frequency defined in Eq. (23). To estimate the neoclassical transport predicted by JONTA we will consider a large tokamak device at low density. Specifically, we will take the density to be $n_{i}=10^{18}\;[1\text{/m}^{3}]$, the minor radius $a=2\;[\text{m}]$, a small inverse aspect ratio of $a/R_{0}=1/6$, and a large toroidal magnetic field of $B_{0}=10\;[\text{T}]$. By considering a low density plasma this ensures the system is in the banana transport regime throughout the majority of the device, whereas the large device size and strong magnetic field results in the banana orbits of the fast ions being much smaller than the system size, a necessary condition for estimating an average diffusivity. A comparison of the neoclassical diffusivity estimated by JONTA with Eq. (25) is shown as a function of minor radius in Fig. 2. Here, good agreement is evident between the computed neoclassical diffusivity and the approximate analytic expression. In particular, the maximum deviation is roughly 12%, with a typical deviation less than 10%. Note that we have removed the inner 20% of the plasma radius from the comparison, since the de-trapping time scales with $\nu_{D}/\varepsilon$, and is thus very short near the magnetic origin, such that this region is not in the banana transport regime.
@@ III.3 Example Ion Orbits and Coordinate Conventions
Figure 3: Circular flux surface geometry used in this work and example collisionless ion orbits. Panel (a) shows co-current ions with a passing ion with initial pitch $\xi=0.8$ (red curve), and a trapped ion with initial pitch $\xi=0.3$ (green curve). Panel (b) shows counter-current ions with a passing ion with initial pitch $\xi=-0.8$ (red curve), and a trapped ion with initial pitch $\xi=-0.3$ (green curve). Flux surface contours are shown in blue. The black marker indicates the initial position of the ions, and arrows indicate their direction. The grad-B drift points downward in our example geometry. We assumed a deuterium ion with $20\;\text{keV}$, a tokamak with a minor radius of $0.5\;\text{m}$, inverse aspect ratio $a/R_{0}=1/3$, constant $q$-profile of $q=2$, and an on-axis magnetic field strength of $B_{0}=2\;\text{T}$.
@@
Before discussing solutions to the adjoint drift kinetic equation, it will be useful to describe the magnetic geometry employed. A plot of the circular flux surface geometry assumed, together with four example ion orbits, are shown in Fig. 3. Here, the blue contours indicate magnetic flux surfaces, and the red and green curves indicate example passing and trapped ion orbits, respectively. The toroidal magnetic field and current are assumed to come out of the page, with the poloidal magnetic field in the counter-clockwise direction. Passing ions with a positive pitch (co-current), will thus rotate in the counter-clockwise direction when their motion is projected onto the poloidal plane as they propagate along a magnetic field line, with a grad-B drift vertically downward. Ions with a negative pitch (counter-current) will move in a clockwise sense when moving along field lines, but will still have a vertically downward grad-B drift. From Fig. 3, it is apparent that co-current ions [Fig. 3(a)] born on the weak field side will tend to be better confined compared to counter-current ions [Fig. 3(b)] born at the same location, due to the magnetic drifts taking the counter-current ions closer to the spatial boundary. Furthermore, a counter-current passing ion when scattered into the trapped region will move onto an orbit that takes it closer to the plasma edge, as evident from Fig. 3(b), such that these ions will be more susceptible to being lost.
IV Mean Escape Time
@@
Figure 4: Loss histories for 20 keV (panel a.) and 50 keV (panel b.) ions. Solid lines indicate the training loss whereas ‘x’ markers indicate the test loss. Six million training points were used and two million test points. The same seed for the pseudorandom number generator was used in both cases.
IV.1 Training Details of the PINN
@@
This section will describe solutions for the mean escape time of a fast ion population at 20 and 50 keV utilizing the PINN implementation described in Sec. III.1, along with a comparison to solutions evaluated with JONTA. Loss histories for ions with 20 and 50 keV are shown in Fig. 4. For these cases, we assumed a midsize tokamak with a minor radius of $a=0.5\;\text{m}$, aspect ratio $R_{0}/a=3$, and deuterium fast ions. The background deuterium plasma is assumed to have a temperature of $1\;\text{keV}$, a density of $n_{i}=10^{20}\;\text{m}^{-3}$, and the safety factor is taken to be $q=2$. A representative Coulomb logarithm of $\ln\Lambda=15$ was used in the collision operator. Each of these quantities were taken to be constant across the plasma radius, though nontrivial spatial profiles could easily be introduced. A fully connected feedforward network with seven hidden layers each with a width of fifty neurons was used. A network of this size, together with six million training points used the majority of the 180 GB of VRAM present on the NVIDIA Blackwell 200 GPU used to train the network. While we have observed improvements in predictions of the PINN by increasing the size of the network, we found the present network size and number of training points offers a good balance in ensuring dense coverage over the ion phase space, while providing a sufficiently expressive network to approximate the solution. For a network of this size, the inference time on an M3 MacBook Pro was measured to be two microseconds per prediction.
The SOAP optimizer was used for the first 50,000 epochs, with SSBroyden used for the remaining 50,000 epochs. Six million training points were used with two million uniformly distributed test points applied. An adaptive residual based sampling method was employed during the SOAP phase of training Wu et al. (2023), such that training points are redistributed near regions with large residuals. From Fig. 4, it is evident that SOAP drives a slow decay of the loss, with SSBroyden driving a sharper decline. The periodic spikes in the training loss during the SOAP phase are due to the training points being resampled every 8,000 epochs. As a result of training points being concentrated in regions with large residuals, the test loss is consistently lower than the training loss, except for the last 10,000 epochs of training for the case of 20 keV ions. This crossing of the test and training loss is due to a large residual forming at some isolated regions immediately adjacent to $r=a$, and will be discussed further below. We note that the SOAP optimizer being employed in this paper is able to efficiently train on a GPU, whereas the SciPy implementation of SSBroyden that we have employed primarily uses a CPU. As a result, the vast majority of the training time is spent during the SSBroyden phase of optimization.
Figure 5: (a) Mean escape time for an ion initialized in the counter-current direction with a pitch of $\xi=-0.8$, after different stages of training. The profile of the ion’s escape time is shown in the panels (a), (b) and (c), with the residual indicated in panels (d), (e) and (f). The first column is after the SOAP phase of training (50,000 epochs), the second column is 10,000 into the SSBroyden phase (60,000 net epochs), and the final column is after 50,000 epochs of SSBroyden (100,000 epochs total). The quantity plotted is $\text{log}_{10}\left(1+T\right)$, yielding a $\text{log}_{10}$ scale for large values of $T$, but vanishing when $T=0$. Deuterium ions with $20\;\text{keV}$ were assumed.
@@
Considering the evolution of the predicted solution during training, Fig. 5(a) shows the predicted mean escape time for a 20 keV ion after the SOAP optimization phase, 10,000 epochs into the SSBroyden phase [Fig 5(b)], and the final predicted mean escape time [Fig 5(c)]. A reference solution from JONTA is shown in Fig. 10(d) below for comparison. It is evident that SOAP is able to approximately capture the structure of the solution near the tokamak edge, but fails to recover the long mean escape time of ions initially located at small minor radii. This is due to the large discrepancy in time scales between ions that begin on loss orbits, and are thus lost on a transit time scale, and those initially located deep in the plasma interior. These latter ions must collisionally diffuse out of the plasma, and are thus lost on a much longer time scale, roughly $10^{5}$ slower for the present example. After 10,000 epochs with SSBroyden, the training and test loss of the residual and boundary terms drop further, and it is evident from Fig. 5 that the relatively long confinement time of ions initially located in the interior is better captured, though its value is still substantially under predicted. Finally, after 50,000 epochs of SSBroyden, the predicted escape time of ions initially located in the interior slowly increases, though the final value is still below that predicted by the particle-based code JONTA. It has been verified that additional training epochs with SSBroyden leads to a more accurate recovery of the mean escape time of the best confined ion orbits, though as described below, begins to overfit the solution near the plasma edge at isolated locations.
@@
Figure 6: Test loss distribution for 20 keV ions. A uniform random distribution with one million points was used. The size of the markers is proportional to the magnitude of the residual.
@@
Turning to the distribution of training losses, two dimensional projections of the resulting test error are shown in Fig. 6, with the magnitude of the markers set to be proportional to the magnitude of the residual. Hence, darker regions represent regions with the largest residuals. From Fig. 6, the outer thirty percent of the tokamak contained the largest losses. This follows due to this region containing direct loss orbits, whose detailed form can be difficult to resolve. It is also evident from Fig. 6(d), where the magnitude of the residuals are plotted as a function of the minor radius, that for $r\approx a$ a small number of very large residuals are present, with the largest having a magnitude of $0.589$ located at $r/a\approx 0.998$ and $\theta\approx 3.041$. These large residuals at the very edge of the plasma are responsible for the increase in the test loss during the last 10,000 epochs of training, but as discussed further below, do not substantially impact the accuracy of the solution throughout the majority of the plasma volume. Finally, regions with $r/a>0.99$ and poloidal angles on either the outboard ($\Delta\theta<\theta<\Delta\theta$) or inboard ($\theta>-\pi+\Delta\theta$ and $\theta<\pi-\Delta\theta$) midplanes were set to zero, leading to empty regions in Fig. 6(a). For this example we set $\Delta\theta=0.1$. As described further below, these regions contain discontinuities in the solution arising from changes in boundary conditions between the lower and upper half of the tokamak, which cannot be precisely resolved. By removing training points from this region, this allows the PINN to find a solution that smoothly transitions across these narrow regions. We note that the largest residuals in the test loss are located at poloidal angles just outside of these regions, and are thus due to the sharp variation of the solution between the upper and lower regions of the tokamak near the plasma edge [see Fig. 9(a) below].
IV.2 Phase Space Structure of the Mean Escape Time